# ScNeuGM: Scalable Neural Graph Modeling for Coloring-Based Contention and Interference Management in Wi-Fi 7

### v1

@@ III-D1 Implicit Loss on Edge Values
However, the partial derivative of the loss w.r.t. each edge value, $\partial L(\theta^{\mu})/\partial E_{i,j}$, can hardly be explicitly estimated. This is because the graph’s performance is measured either STA-wise (e.g., each STA’s reliability) or network-wide (e.g., the number of slots used), and there is no edge-wise (or STA-pairwise) feedback. Meanwhile, the graph’s performance is highly sensitive to individual edge values. For example, if an STA pair is heavily contending or interfering with each other but the NGM NN fails to set $1$ on the edge between them, the reliability constraints of the STAs could be violated. Conversely, if the NGM NN sets $1$ on edges between non-contending and non-hidden STAs, the chromatic number of the graph will increase, preventing the minimization of the number of slots. Due to the implicit nature of STA-pairwise feedback and the performance’s sensitivity to individual edge values, edge gradient estimation in (15) fails to provide efficient update directions for the NGM NN parameters. To address this issue, ScNeuGM uses the ES algorithm, an edge-feedback-free method, for estimating the update directions of the NGM NN in Section IV.
@@ III-D2 Quadratic Complexity In Network Size
The proof is in the appendix. ∎
The above statement implies that any optimal graph remains optimal if edges between non-contending and non-hidden STA pairs are removed. It further suggests that the NGM only needs to consider the edges for contending and hidden pairs, while the remaining edges can be ignored and set to 00. However, since the indicators of contending and hidden pairs, $O^{\mathrm{C}}_{i,j}$ and $O^{\mathrm{H}}_{i,j}$ $\forall i,j$, are not measured, it is unknown which STA pairs require the NGM operations. To address this issue, ScNeuGM uses a low-complexity method based on the DHF to efficiently select the contending and hidden STA pairs in Section V.
IV Parameter Space Optimization of NGM NN Using Evolution Strategy
This section first presents the structure of the NGM NN, followed by the ES algorithm training the NN in the ScNeuGM.
IV-A Design of NGM NN Structure
The structure of the NGM NN $\mu(\cdot|\theta^{\mu})$ is shown in Fig. 5, consisting of 1) a state embedding that maps the STA state sequences in (3) into fixed-dimensional embedding vectors, 2) predictors that infer contention and interference between STA pairs, and 3) an edge generator that outputs the binary edge value. The design of each component is explained as follows.
Figure 5: The structure of the NGM NN $\mu(\cdot|\theta^{\mu})$. Here, the state embedding with parameters $\theta^{\mu}_{\mathrm{SE}}$, the predictors with parameters $\theta^{\mu}_{\mathrm{PC}}$ and $\theta^{\mu}_{\mathrm{PH}}$ are trained using unsupervised/supervised learning in the appendix, and the edge generator with parameters $\theta^{\mu}_{\mathrm{EG}}$ are trained using the ES in Algorithm 2.
@@ IV-A2 Contention and Interference Predictors
Using STA embedding vectors, we can predict how likely each STA pair is contending or hidden using NNs, which assists in deciding edge values. For instance, we use two fully connected NNs, the PCNN and the PHNN, to predict the indicators of contenting and hidden pairs, $O^{\mathrm{C}}_{i,j}$ and $O^{\mathrm{H}}_{i,j}$, based on the embedding vectors $\mathbf{v}_{i}$ and $\mathbf{v}_{j}$ as
@@ IV-A3 Edge Generator
We extract key information on contention and interference to assist the NGM NN in deciding edge values. Specifically, for a given STA pair $i$ and $j$ (where $i\neq j$), the edge value $E_{i,j}$ is generated based on the following information in $\mathcal{S}_{i}$ and $\mathcal{S}_{j}$, representing the negative impact of STA $i$’s transmissions on those of STA $j$ [35], as
  • $\tilde{O}^{\mathrm{C}}_{i,j}$ and $\tilde{O}^{\mathrm{H}}_{i,j}$: Probabilities of the STA pair being contending or hidden, predicted based on their state embeddings. When determining $E_{i,j}$, edge value generation integrates these probabilities with $s_{i,\hat{a}_{i}}$, $s_{j,\hat{a}_{j}}$, and $s_{i,\hat{a}_{j}}$.
We design the edge generator NN (EGNN) based on the extracted contention and interference information as
$E_{i,j}=\mathrm{EGNN}(s_{i,\hat{a}_{i}},s_{i,\hat{a}_{j}},s_{j,\hat{a}_{j}},\tilde{O}^{\mathrm{C}}_{i,j},\tilde{O}^{\mathrm{H}}_{i,j}|\theta^{\mu}_{\mathrm{EG}}),\ \forall i\neq j.$ (19)
Note that the EGNN is configured to output a value in $[0,1]$, where the edge value is rounded to the nearest integer, 00 or $1$, without explicit notation for simplicity.
In summary, the NGM NN parameters consist of the SENN parameters $\theta^{\mu}_{\mathrm{SE}}$, the predictor parameters $\theta^{\mu}_{\mathrm{PC}}$ and $\theta^{\mu}_{\mathrm{PH}}$, and the edge generator parameters $\theta^{\mu}_{\mathrm{EG}}$. When generating the edge for an STA pair $(i,j)$, $\forall i\neq j$, the NGM NN first embeds the STA states $\mathcal{S}_{i}$ and $\mathcal{S}_{j}$ using the SENN as in (16). It then predicts the contention and interference indicators based on the embeddings using the PCNN and the PHNN as in (18), and finally generates the edge using the EGNN as in (19).
IV-B Evolution Strategy for Training the NGM NN
Since the ES algorithm is a parameter space algorithm, we pre-train some parts of the NGM NN to reduce the number of parameters of the NN to be trained by the ES. Specifically, the state embedding and the predictors are pre-trained using unsupervised learning and supervised learning, respectively, as shown in the appendix. Consequently, these parameters in the NGM NN are fixed after pre-training, while only the parameters in the EGNN are updated by the ES. The detailed implementation of the ES is presented as follows.
Figure 6: The ES algorithm that trains the EGNN in the NGM NN.
IV-B1 Reward Design
@@
$R=\begin{cases}\log(\frac{Z^{*}}{Z}),\ \text{if }r_{k}\geq\hat{r},\ \forall k\ ,\\ \log(\min\{\frac{Z^{*}}{Z},1\}\cdot\frac{1}{K}\sum_{k=1}^{K}\min\{\frac{r_{k}}{\hat{r}},1\}),\ \text{otherwise}.\end{cases}$ (20)
If the reliability constraints are satisfied, the reward focuses on minimizing the number of slots; otherwise, the reward emphasizes satisfying the reliability constraints. The maximum reward signal in (20) is 0, and it is achieved if and only if the slot assignments minimize the number of slots while ensuring STA reliability. Note that the minimum number of slots $Z^{*}$ is unknown. Thus, we use the chromatic number of the contention and interference graph, with edges defined as $(O^{\mathrm{C}}_{i,j}+O^{\mathrm{H}}_{i,j})$ for $i\neq j$, as an approximation of $Z^{*}$. To normalize the reward signal, we measure the average reward over the iterations as $\bar{R}$ and aim to maximize the advantage [42, 45] of the reward with respect to the past average, as $R-\bar{R}$. Therefore, the loss of the EGNN parameters in the NGM is designed as
$L(\theta^{\mu}_{\mathrm{EG}})=-\operatorname{\mathbb{E}}[R-\bar{R}].$ (21)
IV-B2 Parameter Space Update
We optimize the EGNN parameters using a Gaussian ES algorithm [38]. In the ES algorithm, each parameter of the EGNN is assumed to be a Gaussian random variable, independent of each other, with the mean $m^{\mu}_{\mathrm{EG}}$ and the variance $\exp({\nu^{\mu}_{\mathrm{EG}}})$. Here, the exponential ensures the variance is always positive. The ES algorithm adjusts each parameter’s mean and variance to find the optimal values that maximize the reward signal, as shown in Fig. 6. Specifically, it first randomly samples parameters $\tilde{\theta}^{\mu}_{\mathrm{EG}}$ from the current distribution as
@@ IV-B2 Parameter Space Update
Using the sampled EGNN parameters $\tilde{\theta}^{\mu}_{\mathrm{EG}}$, the NGM NN generates the edges in the graph based on the network states $\hat{\mathcal{S}}$, as described in Section IV-A. The instant reward, $\dot{R}$, is then evaluated as (20) from the network based on the corresponding slot assignments by coloring the graph. The gradient of the loss function with respect to the parameterized distribution (the mean and the variance) of the EGNN parameters is
$-\nabla_{m^{\mu}_{\mathrm{EG}},\nu^{\mu}_{\mathrm{EG}}}L(\theta^{\mu}_{\mathrm{EG}})=\nabla_{m^{\mu}_{\mathrm{EG}},\nu^{\mu}_{\mathrm{EG}}}\operatorname{\mathbb{E}}\big{[}R-\bar{R}\big{]}$ (23)
=$\stackrel{{\scriptstyle\text{(a)}}}{{=}}\operatorname{\mathbb{E}}_{\theta^{\mu}_{\mathrm{EG}}\sim\mathcal{N}(m^{\mu}_{\mathrm{EG}},\exp({\nu^{\mu}_{\mathrm{EG}}}))}\big{[}(R-\bar{R})\nabla_{m^{\mu}_{\mathrm{EG}},\nu^{\mu}_{\mathrm{EG}}}\log p(\theta^{\mu}_{\mathrm{EG}})\big{]}$
$\stackrel{{\scriptstyle\text{(b)}}}{{\approx}}(\dot{R}-\bar{R})\nabla_{m^{\mu}_{\mathrm{EG}},\nu^{\mu}_{\mathrm{EG}}}\log p(\tilde{\theta}^{\mu}_{\mathrm{EG}})\ ,$
@@
$-\nabla_{m^{\mu}_{\mathrm{EG}}}L(\theta^{\mu}_{\mathrm{EG}})$ $\approx(\dot{R}-\bar{R})\cdot\frac{\tilde{\theta}^{\mu}_{\mathrm{EG}}-m^{\mu}_{\mathrm{EG}}}{\exp({\nu^{\mu}_{\mathrm{EG}}})};$ (24)
$-\nabla_{\nu^{\mu}_{\mathrm{EG}}}L(\theta^{\mu}_{\mathrm{EG}})$ $\approx(\dot{R}-\bar{R})\cdot(\frac{(\tilde{\theta}^{\mu}_{\mathrm{EG}}-m^{\mu}_{\mathrm{EG}})^{2}}{2\exp({\nu^{\mu}_{\mathrm{EG}}})}-\frac{1}{2}).$
@@
V Accelerated Training and Inference of NGM NN Using DHF
Figure 7: The batching mode and the bucketing model of the DHF, which selects a subset of contending and interfering STAs for the training and the inference of the NGM NN, respectively.
This section explains the DHF in the ScNeuGM that accelerates the NGM by focusing the training and inference of the NGM NN only in those contending and hidden STA pairs. We first explain the design of the DHF in Section V-A and then show the usage of the DHF in the training and the inference of the NGM NN in Sections V-B and V-C, respectively.
@@ V-A Design of Deep Hashing Function
We use a DHF NN with parameters $\theta_{\mathrm{HF}}$ to compute each STA’s hash code based on the STA’s state embeddings as
$\mathbf{b}_{k}\triangleq[b_{k}^{1},\dots,b_{k}^{\Lambda}]=\mathrm{DHF}(\mathbf{v}_{k}|\theta_{\mathrm{HF}}),\ \forall k,$ (25)
where $b_{k}^{n}\in\{-1,1\}$ for $n=1,\dots,\Lambda$, and $\Lambda$ is the number of bits in each hash code. To train the DHF, part of the loss is designed to adjust the similarity between codes based on the contention and interference indicators $O^{\mathrm{C}}_{i,j}+O^{\mathrm{H}}_{i,j}$ as
$L^{\prime}(\theta_{\mathrm{HF}})=\frac{1}{|\mathcal{E}|}\sum_{i\neq j}\big{[}\frac{1}{2\Lambda}(\sum_{n=1}^{\Lambda}b_{i}^{n}b_{j}^{n}+\Lambda)-(O^{\mathrm{C}}_{i,j}+O^{\mathrm{H}}_{i,j})\big{]}^{2},$ (26)
where $\frac{1}{2\Lambda}(\sum_{n=1}^{\Lambda}b_{i}^{n}b_{j}^{n}+\Lambda)$ is the normalized Hamming distance between STAs $i$ and $j$’s hash codes. The loss in (26) reduces the Hamming distances between STAs’ codes if they are contending or hidden from each other, or otherwise increases the Hamming distances. Further, to ensure that each bit encodes different information about the STAs’ states, part of the loss is designed to reduce the correlation between bit positions in the hash codes as [39]
$L^{\prime\prime}(\theta_{\mathrm{HF}})=\frac{1}{K^{2}}\|[\mathbf{b}_{1},\dots,\mathbf{b}_{K}][\mathbf{b}_{1},\dots,\mathbf{b}_{K}]^{\rm T}-\mathbb{I}^{K\times K}\|_{2}^{2},$ (27)
where $[\mathbf{b}_{1},\dots,\mathbf{b}_{K}][\mathbf{b}_{1},\dots,\mathbf{b}_{K}]^{\rm T}$ is the matrix with off-diagonal elements that measure the correlations between bit positions. The total loss of the DHF is
@@
where $\lambda>0$ is the weight of the correlation loss part in the total loss. The training of the DHF is outlined in Algorithm 1. The applications of the DHF are explained below.

### v2

@@ III-D1 Implicit Credit Assignment on Edge Values
However, the partial derivative of the loss w.r.t. each edge value, $\partial L(\theta^{\mu})/\partial E_{i,j}$, can hardly be explicitly estimated. This is because the graph’s performance is measured either STA-wise (e.g., each STA’s reliability) or network-wide (e.g., the number of slots used), and there is no edge-wise (or STA-pairwise) feedback. Meanwhile, the graph’s performance is highly sensitive to individual edge values. For example, if an STA pair is heavily contending or interfering with each other but the IGL NN fails to set $1$ on the edge between them, the reliability constraints of the STAs could be violated. Conversely, if the IGL NN sets $1$ on edges between non-contending and non-hidden STAs, the chromatic number of the graph will increase, preventing the minimization of the number of slots. Since many edges in large networks can cause sensitive performance fluctuations, assigning “credit” on how each edge impacts the network is difficult [41]. Thus, edge gradient estimation in (16) fails to provide efficient update directions for the IGL NN parameters when there is no explicit STA-pairwise feedback to individual edges. For example, in a network with many STAs, changing a single edge may not lead to a noticeable change in the overall loss, while the variation in loss is also not directly attributable to specific edges. Therefore, we cannot approximate the gradient of the loss with respect to the parameters, which leads to the failure the DPG/PG methods, as shown in the simulation results. To address this issue, the IGL uses the ES algorithm, an edge-feedback-free method, for estimating the update directions of the IGL NN in Section IV.
@@ III-D2 Quadratic Complexity in Network Size
See Appendix B. ∎
The above statement implies that any optimal graph remains optimal if edges between non-contending and non-hidden STA pairs are removed. It further suggests that the IGL only needs to consider the edges for contending and hidden pairs, while the remaining edges can be ignored and set to 0. However, since the indicators of contending and hidden pairs, $O^{\mathrm{C}}_{i,j}$ and $O^{\mathrm{H}}_{i,j}$ $\forall i,j$, are not measured in online deployment, it is unknown which STA pairs require the IGL operations. To address this issue, the IGL uses a low-complexity method based on the DHF to efficiently select the contending and hidden STA pairs in Section V.
IV Parameter Space Optimization of IGL NN Using Evolution Strategy in IGL
This section first presents the structure of the IGL NN, followed by the ES algorithm training the NN in the IGL.
IV-A Design of IGL NN Structure
The structure of the IGL NN $\mu(\cdot|\theta^{\mu})$ is shown in Fig. 5, consisting of 1) a state embedding that maps the STA state sequences in (3) into fixed-dimensional embedding vectors, 2) predictors that infer contention and interference between STA pairs, and 3) an edge generator that outputs the binary edge value. The design of each component is explained as follows.
Figure 5: The structure of the IGL NN $\mu(\cdot|\theta^{\mu})$. Here, the state embedding with parameters $\theta^{\mu}_{\mathrm{SE}}$, the predictors with parameters $\theta^{\mu}_{\mathrm{PC}}$ and $\theta^{\mu}_{\mathrm{PH}}$ are trained using unsupervised/supervised learning in Appendix C, and the edge generator with parameters $\theta^{\mu}_{\mathrm{EG}}$ are trained using the ES in Algorithm 2. Note that $s_{i,\hat{a}_{i}}$, $s_{i,\hat{a}_{j}}$, and $s_{j,\hat{a}_{j}}$ are extracted from the state sequences $\mathcal{S}_{i}$ and $\mathcal{S}_{j}$.
@@ IV-A2 Contention and Interference Predictors
Using STA embedding vectors, we can predict how likely each STA pair is contending or hidden using NNs, which assists in deciding edge values. For instance, we use two fully connected NNs, the PCNN and the PHNN, to predict the indicators of contending and hidden pairs, $O^{\mathrm{C}}_{i,j}$ and $O^{\mathrm{H}}_{i,j}$, based on the embedding vectors $\mathbf{v}_{i}$ and $\mathbf{v}_{j}$ as
@@ IV-A3 Edge Generator
We extract key information on contention and interference to assist the IGL NN in deciding edge values. Specifically, for a given STA pair $i$ and $j$ (where $i\neq j$), the edge value $E_{i,j}$ is generated based on the information in $\mathcal{S}_{i}$ and $\mathcal{S}_{j}$, representing the negative impact of STA $i$’s transmissions on those of STA $j$ [37], as discussed in Section III-C. We design the edge generator NN (EGNN) based on the extracted contention and interference information as
$E_{i,j}=\mathrm{EGNN}(s_{i,\hat{a}_{i}},s_{i,\hat{a}_{j}},s_{j,\hat{a}_{j}},\tilde{O}^{\mathrm{C}}_{i,j},\tilde{O}^{\mathrm{H}}_{i,j}|\theta^{\mu}_{\mathrm{EG}}),\ \forall i\neq j,$ (20)
where $\tilde{O}^{\mathrm{C}}_{i,j}$ and $\tilde{O}^{\mathrm{H}}_{i,j}$ are predicted based on their state embeddings as (19). Note that the EGNN is configured to output a value in $[0,1]$, where the edge value is rounded to the nearest integer, 0 or $1$, without explicit notation for simplicity.
In summary, the IGL NN parameters consist of the SENN parameters $\theta^{\mu}_{\mathrm{SE}}$, the predictor parameters $\theta^{\mu}_{\mathrm{PC}}$ and $\theta^{\mu}_{\mathrm{PH}}$, and the edge generator parameters $\theta^{\mu}_{\mathrm{EG}}$. When generating the edge for an STA pair $(i,j)$, $\forall i\neq j$, the IGL NN first embeds the STA states $\mathcal{S}_{i}$ and $\mathcal{S}_{j}$ using the SENN as in (17). It then predicts the contention and interference indicators based on the embeddings using the PCNN and the PHNN as in (19), and finally generates the edge using the EGNN as in (20).
IV-B Evolution Strategy for Training the IGL NN
IV-B1 Overview of ES
The ES iterates in the following steps: 1) generate a new candidate IGL NN by adding Gaussian noise to the current parameters; 2) receive a reward signal given this candidate NN; 2) approximate the update direction of the current parameters and the noise variances according to the reward signal. Since the ES algorithm is a parameter space algorithm, we pre-train some parts of the IGL NN to reduce the number of parameters of the NN to be trained by the ES. Specifically, the state embedding and the predictors are pre-trained using unsupervised learning and supervised learning, respectively, as explained in Appendix C. Consequently, these parameters in the IGL NN are fixed after pre-training, while only the parameters in the EGNN are updated by the ES. The detailed implementation of the ES is presented as follows.
Figure 6: The ES algorithm that trains the EGNN in the IGL NN.
IV-B2 Network-Wise Reward Design
@@
[…]
If the reliability constraints are satisfied in the first case, the reward focuses on minimizing the number of slots; otherwise, when the number of slots used is less than the optimal number of slots, e.g., the NN outputs less edge than the optimal graph, the reward emphasizes satisfying the reliability constraints by guiding the algorithm to add more edges and separate more contending or interfering pairs. Note that in the actual system, we do not need to implement the imaginary system that discards the STAs, and it is only for analysis to derive the reward signal. Ideally, we can construct this reward signal if we know $Z^{*}$. Then, the maximum reward signal in (21) is 0, and it is achieved if and only if the slot assignments minimize the number of slots while ensuring STA reliability. However, note that the minimum number of slots $Z^{*}$ is unknown. Thus, when computing $R$, we use the chromatic number of the contention and interference graph, with edges defined as $(O^{\mathrm{C}}_{i,j}+O^{\mathrm{H}}_{i,j})$ for $i\neq j$, to approximate $Z^{*}$ as
$Z^{*}\!\approx\!\tilde{Z}^{*}\!=\!\chi([E_{i,j}|E_{i,j}=O^{\mathrm{C}}_{i,j}+O^{\mathrm{H}}_{i,j},\forall i\neq j;E_{k,k}=0,\forall k]),$ (22)
where $\chi(\cdot)$ returns the number of colors of the input graph’s adjacency matrix using greedy coloring [45]. Since $Z^{*}\leq\tilde{Z}^{*}$, the approximated reward signal can be greater than 0 and has an unknown maximum at $\log\frac{\tilde{Z}^{*}}{Z^{*}}$ (note that $\log\frac{\tilde{Z}^{*}}{Z^{*}}\geq 0$). For simplicity, we use the same notation $R$ as the approximated reward signal in this paper.
To normalize the reward signal, we measure the average reward over the iterations as $\bar{R}$ and aim to maximize the advantage [45, 48] of the reward with respect to the past average, as $R-\bar{R}$. Therefore, the loss of the EGNN parameters in the IGL is designed as
$L(\theta^{\mu}_{\mathrm{EG}})=-\operatorname{\mathbb{E}}[R-\bar{R}|\theta^{\mu}_{\mathrm{EG}}].$ (23)
IV-B3 Parameter Space Update Direction Estimation
We optimize the EGNN parameters using a Gaussian ES algorithm [40]. In the ES algorithm, each parameter of the EGNN is assumed to be a Gaussian random variable, independent of each other, with the mean $m^{\mu}_{\mathrm{EG}}$ and the variance $\exp({\nu^{\mu}_{\mathrm{EG}}})$. Here, the exponential ensures the variance is always positive. The ES optimizes the Gaussian-smoothed surrogate of the loss function as [40]
$L(\theta^{\mu}_{\mathrm{EG}})\approx\operatorname{\mathbb{E}}_{\tilde{\theta}^{\mu}_{\mathrm{EG}}\sim\mathcal{N}(m^{\mu}_{\mathrm{EG}},\exp({\nu^{\mu}_{\mathrm{EG}}}))}\big[L(\tilde{\theta}^{\mu}_{\mathrm{EG}})\big].$ (24)
The ES algorithm adjusts each parameter’s mean and variance to find the optimal values that maximize the reward signal, as shown in Fig. 6. Specifically, it first randomly samples parameters $\tilde{\theta}^{\mu}_{\mathrm{EG}}$ from the current distribution as
@@ IV-B3 Parameter Space Update Direction Estimation
Using the sampled EGNN parameters $\tilde{\theta}^{\mu}_{\mathrm{EG}}$, the IGL NN generates the edges in the graph based on the network states $\hat{\mathcal{S}}$, as described in Section IV-A. The instant reward, $R$, is then evaluated as (21) from the network based on the corresponding slot assignments by coloring the graph. The gradient of the loss function with respect to the parameterized distribution (the mean and the variance) of the EGNN parameters based on (24) is
$\ \ \ -\nabla_{m^{\mu}_{\mathrm{EG}},\nu^{\mu}_{\mathrm{EG}}}L(\theta^{\mu}_{\mathrm{EG}})$ (26)
$\approx-\nabla_{m^{\mu}_{\mathrm{EG}},\nu^{\mu}_{\mathrm{EG}}}\operatorname{\mathbb{E}}_{\tilde{\theta}^{\mu}_{\mathrm{EG}}\sim\mathcal{N}(m^{\mu}_{\mathrm{EG}},\exp({\nu^{\mu}_{\mathrm{EG}}}))}\big[L(\tilde{\theta}^{\mu}_{\mathrm{EG}})\big]$
$=\nabla_{m^{\mu}_{\mathrm{EG}},\nu^{\mu}_{\mathrm{EG}}}\operatorname{\mathbb{E}}_{\tilde{\theta}^{\mu}_{\mathrm{EG}}\sim\mathcal{N}(m^{\mu}_{\mathrm{EG}},\exp({\nu^{\mu}_{\mathrm{EG}}}))}\big[\operatorname{\mathbb{E}}[R-\bar{R}|\theta^{\mu}_{\mathrm{EG}}]\big]$
$\stackrel{{\scriptstyle\text{(a)}}}{{=}}\operatorname{\mathbb{E}}_{\tilde{\theta}^{\mu}_{\mathrm{EG}}\sim\mathcal{N}(m^{\mu}_{\mathrm{EG}},\exp({\nu^{\mu}_{\mathrm{EG}}}))}\big[\operatorname{\mathbb{E}}[R-\bar{R}|\theta^{\mu}_{\mathrm{EG}}]\cdot$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\nabla_{m^{\mu}_{\mathrm{EG}},\nu^{\mu}_{\mathrm{EG}}}\log p(\tilde{\theta}^{\mu}_{\mathrm{EG}})\big]$
$\stackrel{{\scriptstyle\text{(b)}}}{{\approx}}(R-\bar{R})\nabla_{m^{\mu}_{\mathrm{EG}},\nu^{\mu}_{\mathrm{EG}}}\log p(\tilde{\theta}^{\mu}_{\mathrm{EG}}),$
@@
$-\nabla_{m^{\mu}_{\mathrm{EG}}}L(\theta^{\mu}_{\mathrm{EG}})$ $\approx(R-\bar{R})\cdot\frac{\tilde{\theta}^{\mu}_{\mathrm{EG}}-m^{\mu}_{\mathrm{EG}}}{\exp({\nu^{\mu}_{\mathrm{EG}}})};$ (27)
$-\nabla_{\nu^{\mu}_{\mathrm{EG}}}L(\theta^{\mu}_{\mathrm{EG}})$ $\approx(R-\bar{R})\cdot(\frac{(\tilde{\theta}^{\mu}_{\mathrm{EG}}-m^{\mu}_{\mathrm{EG}})^{2}}{2\exp({\nu^{\mu}_{\mathrm{EG}}})}-\frac{1}{2}).$
@@
V Accelerated Training and Inference of IGL NN Using DHF in IGL
Figure 7: The batching mode and the bucketing model of the DHF, which selects a subset of contending and interfering STAs for the training and the inference of the IGL NN, respectively.
This section explains the DHF in the IGL that accelerates the IGL by focusing the training and inference of the IGL NN only on those contending and hidden STA pairs. We first explain the design of the DHF in Section V-A and then show the usage of the DHF in the training and the inference of the IGL NN in Sections V-B and V-C, respectively.
@@ V-A Design of Deep Hashing Function
We use a DHF NN to approximate the LSH, where the NN encodes each STA’s state as a code vector. The similarity between the codes of two STAs approximates the geometric proximity and how likely the contention and interference will occur between them. Since the exact location of STAs is not measured, we use the state embeddings of STA pairs as the input of hashing to approximate the geometric relation in each STA pair. Then, by comparing the bits of the STAs’ hash codes, we can efficiently find the contending and interfering STA pairs. Specifically, the DHF NN with parameters $\theta_{\mathrm{HF}}$ to compute each STA’s soft hash code based on the STA’s state embeddings as
$\tilde{\mathbf{b}}_{k}\triangleq[\tilde{b}_{k}^{1},\dots,\tilde{b}_{k}^{\Lambda}]=\mathrm{DHF}(\mathbf{v}_{k}|\theta_{\mathrm{HF}}),\ \forall k,$ (28)
where $\tilde{b}_{k}^{n}\in[-1,1]$ for $n=1,\dots,\Lambda$, and $\Lambda$ is the number of bits in each hash code. To train the DHF, part of the loss is designed to adjust the similarity between codes based on the contention and interference indicators $O^{\mathrm{C}}_{i,j}+O^{\mathrm{H}}_{i,j}$ as
$L^{\prime}(\theta_{\mathrm{HF}})=\frac{1}{|\mathcal{E}|}\sum_{i\neq j}\big[\frac{1}{2\Lambda}(\sum_{n=1}^{\Lambda}\tilde{b}_{i}^{n}\tilde{b}_{j}^{n}+\Lambda)-(O^{\mathrm{C}}_{i,j}+O^{\mathrm{H}}_{i,j})\big]^{2},$ (29)
where $\frac{1}{2\Lambda}(\sum_{n=1}^{\Lambda}\tilde{b}_{i}^{n}\tilde{b}_{j}^{n}+\Lambda)$ is the normalized Hamming distance between STAs $i$ and $j$’s hash codes. The loss in (29) reduces the Hamming distances between STAs’ codes if they are contending or hidden from each other, or otherwise increases the Hamming distances. Further, to ensure that each bit encodes different information about the STAs’ states, part of the loss is designed to reduce the correlation between bit positions in the hash codes as [42]
$L^{\prime\prime}(\theta_{\mathrm{HF}})=\frac{1}{K^{2}}\|[\tilde{\mathbf{b}}_{1},\dots,\tilde{\mathbf{b}}_{K}][\tilde{\mathbf{b}}_{1},\dots,\tilde{\mathbf{b}}_{K}]^{\rm T}-\mathbb{I}^{K\times K}\|_{2}^{2},$ (30)
where $\tilde{\mathbf{b}}_{1},\dots,\tilde{\mathbf{b}}_{K}][\tilde{\mathbf{b}}_{1},\dots,\tilde{\mathbf{b}}_{K}]^{\rm T}$ is the matrix with off-diagonal elements that measure the correlations between bit positions. The total loss of the DHF is
@@
where $\lambda>0$ is the weight of the correlation loss part in the total loss. The training of the DHF is outlined in Algorithm 1. The hard hash code is generated using the trained DHF as
$\mathbf{b}_{k}\triangleq[b_{k}^{1},\dots,b_{k}^{\Lambda}]=[\operatorname{\bf sgn}\tilde{b}_{k}^{1},\dots,\operatorname{\bf sgn}\tilde{b}_{k}^{\Lambda}],\ \forall k,$ (32)
where $\operatorname{\bf sgn}$ takes the sign, 1-1 or $1$, of the soft hard code bits. The applications of the trained DHF are explained below.